OpenStudy (sleepyjess):

Please please please help me. Mathematical induction For the given statement Pn, write the statements \(P_1, ~P_k,~ and~ P_{k+1}\) 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (sleepyjess):

@Callisto

ganeshie8 (ganeshie8):

You have \[\large P_n ~: ~2 + 4 + 6 + . . . + 2n = n(n+1) \]

OpenStudy (sleepyjess):

does \(P_1\) just mean substitute 1 for n?

ganeshie8 (ganeshie8):

Yep! \[\large P_1 ~: ~2 = 1(1+1) \]

OpenStudy (sleepyjess):

oh ok :) what about \(P_k\)? I am completely confused on that one :/

ganeshie8 (ganeshie8):

simply replace \(n\) by \(k\) \[\large P_k ~: ~2 + 4 + 6 + . . . + 2k = k(k+1)\]

ganeshie8 (ganeshie8):

\[\large P_{\heartsuit} ~: ~2 + 4 + 6 + . . . + 2\heartsuit = \heartsuit(\heartsuit+1)\]

OpenStudy (sleepyjess):

you make this seem so simple...

ganeshie8 (ganeshie8):

You can guess \(P_{k+1}\)

OpenStudy (sleepyjess):

and for \(P_{k+1}\) would it be \(k+1(k+1+1)\)?

ganeshie8 (ganeshie8):

P_(k+1) is a "statement", not just an expression

ganeshie8 (ganeshie8):

it should say some complete sentence

OpenStudy (sleepyjess):

\(\large P_{k+1} ~: ~2 + 4 + 6 + . . . + 2k = k+1(k+1+1)\)?

ganeshie8 (ganeshie8):

try again, you should replace all "n" by "(k+1)"

OpenStudy (sleepyjess):

\(\large P_{k+1} ~: ~2 + 4 + 6 + . . . + 2k = (k+1)((k+1)+1)\)

ganeshie8 (ganeshie8):

right side looks good, left side isnt..

OpenStudy (sleepyjess):

ugh \(\large P_{k+1} ~: ~2 + 4 + 6 + . . . + 2(k+1) = (k+1)((k+1)+1)\) hopefully?

ganeshie8 (ganeshie8):

Thats it !

OpenStudy (sleepyjess):

:O

OpenStudy (sleepyjess):

yay!

OpenStudy (sleepyjess):

yay!

OpenStudy (sleepyjess):

Thank you so much! I have been stuck on that question for hours

ganeshie8 (ganeshie8):

yw:) principle of induction is pretty interesting topic.. i hope they let you do some induction proofs as well !